OpenStudy (anonymous):

Let Us Cell U charges $0.15/min. for long-distance calls. Cell for Less charges $10.50/hr. plus an additional $0.10/min. for partial hours.If you average 5 hrs. and 15 min. of long distance calls each month, which company will save you money? a.Let Us Cell U b.Cell for Less c.Both are the same d.Not enough information

OpenStudy (anonymous):

Let us cell you would save you a lot of money over Cell for Less charges.

OpenStudy (igreen):

0.15/min = Let Us Cell U Total time: 5hrs, 15 min. 5 hours is 300 minutes, so we have a total of 315 minutes. 315 * 0.15 = 47.25. So Let Us Cell U charges $47.25 10.50/hr + 0.10/min = Cell for Less Total time: 5hrs, 15 min. 5 * 10.5 = 52.5 15 * .10 = 1.5 52.5 + 1.5 = 54 So Cell for Less charges $54.00

OpenStudy (igreen):

Therefore, Let Us Cell U charges LESS and saves you money.

OpenStudy (gorv):

u charge for a hour that is 60 minute =0.15*60=9

OpenStudy (gorv):

so u will save you